There are two faucets in a tub.If one faucet can fill the tub in 10 minutes alone and the other in three minutes, then how long will it take for two faucets to fill the tub if they are both open
Answer by Alan3354(69443)

There are two faucets in a tub.If one faucet can fill the tub in 10 minutes alone and the other in three minutes, then how long will it take for two faucets to fill the tub if they are both open
============
Parallel work/flow/resistance problem.
Look up solved problems.
-------
t = 10*3/(10+3) = 30/13 minutes.